< Previous week | Next week >
IODP Expedition 336:
Mid-Atlantic Ridge Microbiology
Week 6 Report (23-29 October 2011)
PDF file is available for download.
After cementing the 10.75 inch casing in at 60.4 m below seafloor (mbsf), drilling
in Hole U1383C proceeded with an RCB bit from 69.5 to 211.6 mbsf. From
this 142.1 m long interval, 28.55 m of core was recovered (20.1%). Rocks
are glassy to fine-grained basalts with variable phenocryst
(plagioclase and olivine) contents. Down to 127 mbsf, core consists of microcrystalline to fine-grained, sparsely plagioclase phyric basalt
with abundant glassy margins and numerous intervals of hard interflow
limestone. From 127 to 164 mbsf, massive, plagioclase-olivine phyric basalts are developed, which occasionally host limestone (with and without basalt clasts) as fracture fill. Below 164 mbsf, glassy to variolitic to
cryptocrystalline basalts (most likely pillow flows) predominate and limestone
is largely missing. Each of these three main lithologic
units is divided into numerous (between 22 and 48) subunits based on the
occurrence of glassy margins. The overall abundance of glass is noticeably
greater than in Hole U1382A, and the extent of palagonitization
ranges from weak to moderate. Basalts are avesicular
to sparsely vesicular and show vesicle fills of clay, carbonate, and Fe-oxyhydroxide. Brownish alteration halos commonly
track veins filled with clay and/or carbonate and zeolite. After changing
the RCB bit and repairing the heave compensation system, coring resumed in the evening hours of the 29th.
We collected 51 hard rock whole rounds for microbiological
analysis. Samples were selected in
the core splitting room as quickly as possible after core recovery, following
initial discussion with petrologists on sample representation and photographic
documentation of the sample before removal from the core liner. When sample volume permitted, samples
were preserved for shipboard deep UV scanning, shore-based DNA and RNA
analysis, shore-based fluorescence in situ hybridization and cell counting
analysis, shipboard culturing and enrichment, shore-based isotopic analysis,
and shipboard fluorescent microsphere analysis. At least one, and usually several, microbiological hard rock samples were collected from every core section. Hard rock samples span a range of
lithological units, alteration states, presence of chilled margins, and some
contain at least one vein/fracture. Additionally, a few recovered plastic bags
that held the fluorescent microsphere solutions in the core catcher have been
collected as a contamination check in DNA analysis. Examination of the
microsphere abundance in/on the recovered samples is ongoing.
Hole U1383C was initiated last week by (1) installing a re-entry cone
and 34.7 m of 16-inch casing and (2) preparing a hole to install 10.75 inch
casing by drilling a 14.75-inch hole to 69.5 mbsf (31.2 m below the
This week we
continued with operations in Hole U1383C. After completing the drilling and
hole conditioning, we retrieve the drill string and the bit was back on board
at 0702 hours on 23 October. We assembled 60.41 m of 10.75-inch casing and
lowered it to just above the seafloor, pausing about halfway to deploy the camera
system. After ~30 minutes of maneuvering, we reentered Hole U1383C at 1626 hours
on 23 October. The casing string was landed at 1720 hours and we confirmed it
was latched in with a 20,000 lb overpull.
The casing was secured in place by pumping 25 barrels of cement with lost
circulation material (Cello-Flake). The casing running tool was released at
1940 hours, the drill string flushed clear of any remaining cement, and we started pulling out of the hole. With the casing running tool at 2336 mbrf, the trip was halted so we could slip and cut 115' of
drill line. We also spent two hours replacing a spool valve and repairing a
hydraulic line on the 5 inch pipe racker. We then continued retrieving the casing running tool, which was back on board at 0825 hours on 24 October.
We assembled an
RCB and bottom-hole assembly, verified the core barrel space out, and lowered
this to the seafloor. About halfway down, we paused to deploy the camera system.
This reentry only took us ~10 minutes of maneuvering. After we retrieved the
camera system, installed the top drive, and spaced out for drilling, we tagged the
top of the cement at 43.8 mbsf (~16.6 m above the casing shoe). Once we drilled
out the cement and cleaned out the rat hole, we recovered the center bit. We
dropped an RCB core barrel and started coring at 0000 hours on 25 October. We
cut Cores U1383C-2R through -18R from 69.5 to 211.6 mbsf. After Core 15R, we
performed a wiper trip back up to the casing shoe and encountered 20 m of fill
getting back to the bottom of the hole. Hole cleaning remained a priority with
mud sweeps being performed on average twice during each cored interval. A total
of 142.1 m were cored and 28.55 m recovered (20%). Just after starting to cut
Core 18R and with 45.8 hours on the bit, the vessel began experiencing high
heave. This heave made it almost impossible to keep sufficient weight on bit to
keep the bit on bottom, so we circulated cuttings out of the hole with a final
mud sweep and pulled the bit out of the hole. This first RCB bit was back on
board at 1720 hours on 28 October. It was fortunate that we did not continue
coring with this bit as it had experienced bearing failures on all four roller cones.
assembled a new RCB bit (C-7) bit and inspected the float valve and support
bearing assemblies, we conducted routine rig maintenance. This identified a
faulty air cylinder for the locking pin on the motion compensator that had to
be repaired. This resulted in 9.25 hours of rig down time. During the repair,
we assembled the previous bottom-hole assembly and added 3 more drill collars.
We installed a center bit so that the float valve would remain open during the trip allowing the drill string to fill with seawater. At 0630 on 29 October, we
started tripping the BHA to the seafloor. The reentry cone was visible as soon
as the camera system reached the seafloor and the bit was almost directly over
the cone. We reentered Hole U1383C at 1231 on 29 October. We retrieved the
camera system and the center bit and lowered the bit to the bottom of the hole
while carefully checking the hole for tight spots and fill.
Education and Outreach
Outreach efforts for this expedition have continued through a variety of programs.
Blogs: Current bloggers
this week include our onboard education officer Jennifer Magnusson (personal,
educator ideas, and a blog for kids) and staff scientist Adam Klaus
(operations). Heath Mills continues writing about Mid-Atlantic Ridge
microbiology on Texas A&M's georesearch page.
Katrina Edwards continues her blog on the Scientific American Expeditions page
and the C-DEBI site. Beth Orcutt continues to blog
about microbiology on the Adopt-a-Microbe website and Amanda Haddad continues
to provide science content and connect with a special needs audience on the
Classroom Connections website.
Videoconferences: Six live ship-to-shore interactive programs were
conducted with the following audiences:
- 7th/8th graders in California
- 6th – 8th graders in Illinois
- 7- and 8-year-olds in Spain
- 11th/12th graders in Pennsylvania
- Texas A&M University weekly Geology and Geophysics seminar
Ten interactive events are scheduled for next week.
Social Media: The education officer continues to
post daily updates on the JR Facebook page and Twitter account. Updates include
links to the blog or other pages on the JR website, photos, videos, operational
updates, and classroom activities. Daily math questions continued to be posted.
An activity was posted that encourages students to combine math and art to
create a model of a re-entry cone and calculate drilling distances to scale.
Adopt-A-Microbe: Week 6 activities (classes created fabric microbe models) were submitted and Week 7 activities (culturing microbes) were assigned. Microbioloigst Steffen Leth Jorgensen provided an introduction to Methanopyrus kandleri.
Classroom Connection: This week's theme was "Life on board the
JR," and students participated in a variety of related activities. These
included designing and building a kite, watching videos from the JR website,
and an interview with the image specialist, Bill Crawford.
Documentary: The videographers have continued full-time filming and
interviewing for their documentary.
Technical Support and HSE Activities
Science Mission Support: Technical staff provided analytical support for coring operations at Hole U1383C and worked on various maintenance
issues around the labs.
Other Technical Activities:
- Continued to assist scientists with DESCLogik;
- Issues with consistent velocity calibrations were finally resolved by changing parameters in the software code;
- Developers continued to work with staff to correct and test upgrades to Sample Master;
- Work continued on 3D camera project: (a) temporary mounting hardware installed for testing system on the section half multisensory logger, (b) coding and testing of software modules in progress
The weekly fire and abandon ship drill was held as scheduled.